OpenStudy (anonymous):

The surface area of a rectangular prism is found using the formula SA=2(lw+lh+wh), where l is length, w is width, and h is height. If the surface area of a rectangular prism is 324m^2, with length 5m and width 12m. What is the height, in meters? A. 60 B. 5 C. 6 D. -2

OpenStudy (mef4.0):

Sub in what you know from the info they give you, then solve for h: 324 = 2((5*12) + 5h + 12h) 324 = 2(60 + 17h) Then distribute, get h by itself and solve. that make sense?

OpenStudy (anonymous):

so what is the answer? a,b,c,d..